In a tv show Darna, a stuntwoman drives horizontally with a speed of 1.34 m/s off a platform that is 12.0 m above the water. What is her speed just before striking the water?
1
Expert's answer
2020-10-22T18:07:23-0400

In order to safe the conservation law:

"v' = (v^2 + 2gh)^\frac{1}{2} = ((1,34)^2 + 2\times9,8\times1,2)^{\frac{1}{2}} = 5.03"